# thanks: Equal contributionthanks: Equal contribution

### v2

@@ Abstract
Now, we perform a theoretical analysis to explain the observed RT dynamics in our system and to provide a rationale for the $\alpha-D_{r}$ phase diagram. As observed in experiments and numerical simulations, the dynamics of the system at any given time are entirely governed by its internal configuration, which is represented by variables $\theta_{i}$ ($i$ = 1 and 2 for two robots) [Fig. 4(a)]. Therefore, we define another set of generalized angular coordinates: $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$ (see Fig. 4(a)). Note that $\beta$ equals the principal value of the angle $2\theta_{-}$ in the range $[-\pi,\pi]$ and $\theta_{+}$ indicates the average orientation of the robots with respect to the rod. Fig. 4(a) clearly shows that motility causes increasing and decreasing trends in $\theta_{1}$ and $\theta_{2}$, respectively. Therefore, by definition, $\theta_{-}$ also increases with time over large time scales. Moreover, for integers $n$, when $\theta_{-}=n\pi$, the robots are parallel to each other and thus move with the maximum possible speed of $v_{a}$. Similarly, when $\theta_{-}=(2n+1)\pi/2$, the system shows tumbling behavior. To elaborate further, we derive the equations of motion for $\theta_{\pm}$ from Eqs. (30) and (32) of SM, which are given as follows (also see SM section IIC):
$\dfrac{d\theta_{\pm}}{dt}=\mathcal{T}_{\pm}+\sqrt{2D_{\pm}}\eta_{\pm}(t),\$ (1)
@@
where $\mathcal{T}_{+}=(2v_{a}/\ell)\sin\theta_{-}\left(\sin\theta_{+}-\ell\cos\theta_{+}\mathcal{G}\mathcal{H}\right)$, $\mathcal{T}_{-}=-(2\delta v_{a}/\Gamma_{\tau})\mathcal{G}\cos^{2}\theta_{+}\cos(\theta_{-}+\alpha)\sin\theta_{-}$, and $D_{\pm}$, $\mathcal{G}$ and $\mathcal{H}$ are the functions $\theta_{+}$ , $\theta_{-}$, and $\alpha$ (see SM section IIC). The strengths of the noise terms $D_{\pm}$ in Eq. (1) are of the order of $D_{r}$. $\eta_{\pm}(t)$ are the delta-correlated noise with zero mean and variance one.
@@
Figure 4: (a) A schematic diagram clarifying the angular coordinates $(\theta_{+},\theta_{-})$. Here, $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$. (b) Flow diagram of $(\theta_{+},\theta_{-})$ for $\alpha=120^{\circ}$ evaluated from the deterministic part $\mathcal{T}_{\pm}$ of Eq.  (1). $S_{1}$\&& $S_{2}$ and $U_{1}$\&& $U_{2}$ encircled by dotted lines represent stable and untable points respectively. Black dashed lines represent experimental trajectories for $D_{r}=0$. Schematic configurations of $S_{1}$ and $S_{2}$ are shown at the bottom. $\mathbf{n_{1}}$ and $\mathbf{n_{2}}$ are misaligned by an angle of $2\alpha-\pi=60^{\circ}$ in both cases. (c) Flow diagram for $\alpha=90^{\circ}$. Blue dotted lines at $\theta_{-}=n\pi$ are semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$, where $n$ denotes positive integers. Black dashed lines are the experimental trajectories of $D_{r}=0$. (d) Flow profile for $\alpha=60^{\circ}$. The system exhibits four unstable points ($U_{1}$, $U_{2}$, $U_{3}$, and $U_{4}$). Solid blue lines at $\theta_{-}=n\pi$ represent configurations that are stable in $\theta_{-}$ and neutral in $\theta_{+}$. The black dashed experimental trajectories for the $D_{r}=0$ case. Three among infinitely many run configurations corresponding to $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$ are shown at the bottom. The color bars in (b), (c) &\&& (d) correspond to the magnitude of the vector $(\mathcal{T}_{+},\mathcal{T}_{-})$ in rad s-1.
@@
We begin with the case where $\alpha>90^{\circ}$, with a typical flow profile shown for $\alpha=120^{\circ}$ in Fig. 4(b). Here, we find two stable ($S_{1}$, $S_{2}$) and two unstable ($U_{1}$, $U_{2}$) fixed points. We also observe an unstable fixed line along $\theta_{-}=n\pi$ which is less relevant to the results presented. Locations of stable points in the first and second quadrants are $(\theta_{+},\theta_{-})$ = $(\pi,3\pi/2-\alpha)$ and $(0,3\pi/2-\alpha)$ for $S_{1}$ and $S_{2}$ respectively, with system configurations shown below Fig. 4(b). Both configurations correspond to run states. Therefore, the system favors these stable points for low noise values, resulting in a pronounced run regime at low $D_{r}$ values, explaining the experimental results presented in Figs. 3(c) and (e). Interestingly, both $S_{1}$\&& $S_{2}$ correspond to non-aligned robots ($\mathbf{n_{1}}\nparallel\mathbf{n_{2}}$). To test this, we conduct experiments for the $D_{r}=0$ case using random initial conditions in $(\theta_{+},\theta_{-})$ and find that the system follows the predicted flow lines (black dashed lines in Fig. 4(b)) before eventually settling into one of the stable regions, indicated by the dashed circular regions (see SM movie 6). Moreover, as evident from configurations $S_{1}$&\&& $S_{2}$, the system’s run speed at the stable points is given by $v_{a}\sin\alpha$, which decreases with $\alpha$ for $\alpha>90^{\circ}$. This explains the decreasing trend in $V_{\text{run}}$ shown in insets of Figs. 3(e) and 3(f). We also observe that at low but finite $D_{r}$ values, the system runs corresponding to $\theta_{-}$ = $\pi/2-\alpha+n\pi$, which equals $\theta_{-}=n\pi-\pi/6$ for $\alpha=120^{\circ}$ while occasionally transitioning between $S_{1}$ and $S_{2}$ leading to tumbling events [See SM Section IID].
@@
For $\alpha=90^{\circ}$, no stable fixed points are observed. Instead, there exists a fixed line [blue dashed line in Fig. 4(c)] at $\theta_{-}=n\pi$, which is semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$ Therefore, as the system reaches this line, it runs for a while before sliding towards increasing $\theta_{-}$ due to noise. Experiments performed at $D_{r}=0$, in this case, show traced paths that agree well with theoretical values (see SM movie 7). Finally, for $\alpha<90^{\circ}$, and using $\alpha=60^{\circ}$ as a typical example, we find four unstable points ($U_{1}$, $U_{2}$, $U_{3}$, $U_{4}$) and a fixed line at $\theta_{-}=n\pi$ as shown in Fig. 4(d). This fixed line is neutral in $\theta_{+}$ while stable in $\theta_{-}$ corresponding to infinitely possible configurations with $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$ with schematics of three typical configurations shown below the Fig. 4(d). Consequently, $V_{\text{run}}\approx v_{a}$ for $\alpha\leq 90^{\circ}$ as observed experimentally in inset of Fig. 3(e). Again, the grey dashed line indicates the experimental trajectory for $D_{r}=0$ (SM movie 8).
To conclude, inspired by microorganism motility, we present a robotic model system that replicates run-and-tumble-like (RT-like) motion. Our robots, driven by motorized wheels through mechanical gears and responding to microcontroller signals, undergo rolling-without-slipping motion. As a result, they faithfully emulate overdamped active motion, a hallmark of microorganisms operating at extremely low Reynolds numbers. We program both robots to exhibit an overdamped active Brownian (AB) motion. The rigid rod, attached to a pivot point on the robot’s body, rotates freely and facilitates rotational motion. By varying the pivot placement relative to the robot’s polarity axis, we observe a range of rich, tunable RT-like dynamics. We quantify the RT-like motion using the tumbling frequency and identify two key tuning parameters. The first, $\delta$, is the distance of the pivot from the robot’s centre, which promotes tumbling frequency by amplifying torque. The second, $\alpha$, is the angle between the pivot-to-center line and the polarity axis, which also adjusts the tumbling frequency based on rotational noise programmed inside each robot. The tunability of RT motion with these parameters might have implications in the critical role played by inter-flagella connection through the organism’s body in achieving synchronization cycles [23, 25]. We also developed a theoretical model that captures the essential features of our system and reproduces the experimental results, elucidating the intricate dependence of RT dynamics on $\delta$ and $\alpha$.

### v4

@@ Appendix A End Matter
Here, we perform a theoretical analysis to explain the observed RT dynamics in our system and to provide a rationale for the $\alpha-D_{r}$ phase diagram. As observed in experiments and numerical simulations, the dynamics of the system at any given time are entirely governed by its internal configuration, which is represented by variables $\theta_{i}$ (ii = 1 and 2 for two robots) [Fig. 4(a)]. Therefore, we define another set of generalized angular coordinates: $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$ (see Fig. 4(a)). Note that $\beta$ equals the principal value of the angle $2\theta_{-}$ in the range $\pi,\pi]$ and $\theta_{+}$ indicates the average orientation of the robots with respect to the rod. Fig. 4(a) clearly shows that motility causes increasing and decreasing trends in $\theta_{1}$ and $\theta_{2}$, respectively. Therefore, by definition, $\theta_{-}$ also increases with time over large time scales. Moreover, for integers $n$, when $\theta_{-}=n\pi$, the robots are parallel to each other and thus move with the maximum possible speed of $v_{a}$. Similarly, when $\theta_{-}=(2n+1)\pi/2$, the system shows tumbling behavior. To elaborate further, we derive the equations of motion for $\theta_{\pm}$ from Eqs. (30) and (32) of SM [44], which are given as follows (also see SM section IID [44]):
$\dfrac{d\theta_{\pm}}{dt}=\mathcal{T}_{\pm}+\sqrt{2D_{\pm}}\eta_{\pm}(t),\\$ (1)
@@
where $\mathcal{T}_{+}=(2v_{a}/\ell)\sin\theta_{-}\left(\sin\theta_{+}-\ell\cos\theta_{+}\mathcal{G}\mathcal{H}\right)$, $\mathcal{T}_{-}=-(2\delta v_{a}/\Gamma_{\tau})\mathcal{G}\cos^{2}\theta_{+}\cos(\theta_{-}+\alpha)\sin\theta_{-}$, and $D_{\pm}$, $\mathcal{G}$ and $\mathcal{H}$ are the functions $\theta_{+}$ , $\theta_{-}$, and $\alpha$ (see SM section IID [44]). The strengths of the noise terms $D_{\pm}$ in Eq. (1) are of the order of $D_{r}$. $\eta_{\pm}(t)$ are the delta-correlated noise with zero mean and variance one.
@@
Figure 4: (a) A schematic diagram clarifying the angular coordinates $(\theta_{+},\theta_{-})$. Here, $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$. (b) Flow diagram of $(\theta_{+},\theta_{-})$ for $\alpha=120^{\circ}$ evaluated from the deterministic part $\mathcal{T}_{\pm}$ of Eq.  (1). $S_{1}$ &\& $S_{2}$ and $U_{1}$ &\& $U_{2}$ encircled by dotted lines represent stable and unstable points, respectively. Black dashed lines represent experimental trajectories for $D_{r}=0$. Schematic configurations of $S_{1}$ and $S_{2}$ are shown at the bottom. Here, $\mathbf{n_{1}}$ and $\mathbf{n_{2}}$ are misaligned by an angle of $2\alpha-\pi=60^{\circ}$ in both cases. (c) Flow diagram for $\alpha=90^{\circ}$. Blue dotted lines at $\theta_{-}=n\pi$ are semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$, where $n$ denotes positive integers. Black dashed lines are the experimental trajectories for $D_{r}=0$. (d) Flow profile for $\alpha=60^{\circ}$. The system exhibits four unstable points ($U_{1}$, $U_{2}$, $U_{3}$, and $U_{4}$). Solid blue lines at $\theta_{-}=n\pi$ represent configurations that are stable in $\theta_{-}$ and neutral in $\theta_{+}$. The black dashed experimental trajectories for the $D_{r}=0$ case. Three among infinitely many run configurations corresponding to $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$ are shown at the bottom. The color bars in (b), (c) &\& (d) correspond to the magnitude of the vector $(\mathcal{T}_{+},\mathcal{T}_{-})$ in rad s-1.
@@
We begin with the case where $\alpha>90^{\circ}$, with a typical flow profile shown for $\alpha=120^{\circ}$ in Fig. 4(b). Here, we find two stable ($S_{1}$, $S_{2}$) and two unstable ($U_{1}$, $U_{2}$) fixed points. We also observe an unstable fixed line along $\theta_{-}=n\pi$ which is less relevant to the results presented. Locations of stable points in the first and second quadrants are $(\theta_{+},\theta_{-})$ = $(\pi,3\pi/2-\alpha)$ and $(0,3\pi/2-\alpha)$ for $S_{1}$ and $S_{2}$ respectively, with system configurations shown below Fig. 4(b). Both configurations correspond to run states. Therefore, the system favors these stable points for low noise values, resulting in a pronounced run regime at low $D_{r}$ values, explaining the experimental results presented in Figs. 3(c) and (e). Interestingly, both $S_{1}$ &\& $S_{2}$ correspond to non-aligned robots ($\mathbf{n_{1}}\nparallel\mathbf{n_{2}}$). To test this, we conduct experiments for the $D_{r}=0$ case using random initial conditions in $(\theta_{+},\theta_{-})$ and find that the system follows the predicted flow lines (black dashed lines in Fig. 4(b)) before eventually settling into one of the stable regions, indicated by the dashed circular regions (see SM movie 7 [44]). Moreover, as evident from configurations $S_{1}\&S_{2}$, the system’s run speed at the stable points is given by $v_{a}\sin\alpha$, which decreases with $\alpha$ for $\alpha>90^{\circ}$. This explains the decreasing trend in $V_{\text{run}}$ shown in the insets of Figs. 3(e) and 3(f). We also observe that at low but finite $D_{r}$ values, the system runs corresponding to $\theta_{-}$ = $\pi/2-\alpha+n\pi$, which equals $\theta_{-}=n\pi-\pi/6$ for $\alpha=120^{\circ}$ while occasionally transitioning between $S_{1}$ and $S_{2}$ leading to tumbling events [See SM Section IIE [44]].
@@
For $\alpha=90^{\circ}$, no stable fixed points are observed. Instead, there exists a fixed line [blue dashed line in Fig. 4(c)] at $\theta_{-}=n\pi$, which is semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$. Therefore, as the system reaches this line, it runs for a while before sliding towards increasing $\theta_{-}$ due to noise. Experiments performed at $D_{r}=0$, in this case, show traced paths that agree well with theoretical values (see SM movie 8 [44]). Finally, for $\alpha<90^{\circ}$, and using $\alpha=60^{\circ}$ as a typical example, we find four unstable points ($U_{1}$, $U_{2}$, $U_{3}$, $U_{4}$) and a fixed line at $\theta_{-}=n\pi$ as shown in Fig. 4(d). This fixed line is neutral in $\theta_{+}$ while stable in $\theta_{-}$, corresponding to infinitely possible configurations with $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$, with schematics of three typical configurations shown below in Fig. 4(d). Consequently, $V_{\text{run}}\approx v_{a}$ for $\alpha\leq 90^{\circ}$ as observed experimentally in inset of Fig. 3(e). Again, the grey dashed line indicates the experimental trajectory for $D_{r}=0$ (SM movie 9 [44]).